# Social Bias in Popular Question-Answering Benchmarks

### v2

@@ 6 Conclusion
Our work finds significant limitations regarding transparency and social representativeness in 30 popular QA and RC benchmarks. Many of these benchmarks lack information about annotator demographics, recruitment criteria, and language specificity. Many are, furthermore, biased in terms of gender, occupation, religion, and geographic representation. This has objectionable epistemological and ethical implications, e.g., by incentivizing the development of technologies that serve the needs of a privileged few. We highlight the need for rigorous documentation, validation, and representation standards in LLM benchmarking.
@@
Due to the lack of transparency across benchmarks, we were unable to investigate the causal relationship between the identity of those involved in the benchmark creation and the biases found in the QA and RC benchmark datasets through statistical testing. Due to the immense annotation efforts involved, our analyses were limited in scope. Future work should include a larger number and wider range of benchmarks to allow for more generalizable conclusions.
@@ Acknowledgments
This work was mainly conducted during the first author’s Research Fellowship at Weizenbaum Institute, Berlin. This work was also supported by the German Research Foundation (DFG) project NFDI4DS under Grant No.: 460234259 and a NVIDIA Academic Hardware Grant.
@@ Appendix A Full Benchmark Paper Checklist
Domain Benchmark Language
@@
Encyclopedic QuAC -
DROP -
XQuAD - -
SQuAD - -
HotpotQA - - -
StrategyQA - - -
COQA - - -
NaturalQuestions - - - -
TriviaQA - - - -
BoolQ - - - -
WebQuestions - - - -
Commonsense COPA - - -
WinoGrande -
PIQA - - - -
CommonsenseQA - - -
TruthfulQA - - -
HellaSwag - - -
SIQA - - - -
Scholarly BioASQ-QA - -
GPQA
RACE - -
OpenBookQA - -
MATH - - -
ScienceQA - - - -
MMLU - - - - -
GSM8K - - - -
ARC - - - -
Multimodal MMMU - -
TextVQA - - - - -
OK-VQA - - - - -
@@ Appendix C Benchmark Dataset Analysis Ext’d
Table 6: Detailed list of the numbers of Wikidata entities and associated properties extracted for each benchmark.

### v3

@@ 6 Conclusion
Our work finds significant limitations regarding transparency and social representativeness in 30 popular QA and RC benchmarks. Many of these benchmarks lack information about annotator demographics, recruitment criteria, and language specificity. Many are linguistically biased and tend to exhibit biases towards entities of certain gender, occupations, religions, and locations. This has objectionable epistemological and ethical implications, e.g., by incentivizing the development of technologies that serve the needs of a privileged few. We highlight the need for rigorous documentation, validation, and representation standards in LLM benchmarking.
@@
Due to the lack of transparency across benchmarks, we were unable to investigate the causal relationship between the identity of those involved in the benchmark creation and the biases found in the benchmark datasets through statistical testing.
There is a certain risk that the biases of Wikidata and the entity linker may influence our results. This is hard to avoid in an analysis that utilizes automated processes. Especially for the commonsense and scholarly benchmarks, this is to be considered as a limitation. As for the encyclopedic benchmarks, we assume that this to be less of an issue, because many of them are built on top of Wikidata or Wikipedia (which are content-wise very alike), to begin with.
Some time has gone by between conducting this research and the publication of this article. So, it is likely that newer benchmarks would now fall into our selection criteria that we did not consider. Moreover, due to the large annotation efforts required in this study, we had to limit the scope. Therefore, we set strict selection criteria, which happened to exclude multilingual benchmarks. Future work should include a larger number and wider range of benchmarks to allow for more generalizable conclusions. Studies conducted at larger scale should also systematically examine whether benchmarks have become less biased and more transparent over time.
@@ Acknowledgments
This work was funded through a Research Fellowship at Weizenbaum Institute, Berlin. It was also supported by the German Research Foundation (DFG) project NFDI4DS under Grant No.: 460234259 and an NVIDIA Academic Hardware Grant.
@@ Appendix A Full Benchmark Paper Checklist
Year Benchmark Language
@@
Encyclopedic
2018 QuAC -
2019 DROP -
2020 XQuAD - -
2016 SQuAD - -
2018 HotpotQA - - -
2021 StrategyQA - - -
2019 COQA - - -
2019 NaturalQuestions - - - -
2017 TriviaQA - - - -
2019 BoolQ - - - -
2023 WebQuestions - - - -
Commonsense
2012 COPA - -
2021 WinoGrande -
2020 PIQA - - - -
2019 CommonsenseQA - - -
2022 TruthfulQA - - -
2019 HellaSwag - - -
2019 SIQA - - - - -
Scholarly
2023 BioASQ-QA - -
2023 GPQA
2017 RACE - -
2018 OpenBookQA - -
2021 MATH - - -
2022 ScienceQA - - - -
2021 MMLU - - - -
2021 GSM8K - - - - -
2018 ARC - - - -
Multimodal
2024 MMMU - -
2019 TextVQA - - - - -
2019 OK-VQA - - - - -
@@ Appendix C Benchmark Dataset Analysis Ext’d
Table 6: Detailed list of the numbers of Wikidata entities and associated properties extracted for each benchmark. Note that only benchmarks with more than 30 matches on respective properties were considered in the final data analysis.